Benzoic acid is a weak acid that has antimicrobial properties. Its sodium salt, sodium benzoate, is a preservative found in foods, medications, and personal hygiene products. Benzoic acid ionizes in water: C6H5COOH⇌C6H5COO−+H+ The pKa of this reaction is 4.2. In a 0.62 M solution of benzoic acid, what percentage of the molecules are ionized? Express your answer to two significant figures and include the appropriate units.
Answer #1

pka_4:2 log ka=4.2 10 104.2 - ka Ka = 0.000063 = 6,3 x 10 -00+ @ @ 5-04 = +C a it ou -ca 9 I j C-ca e harl d=ionizatia c= conanswer = 1%